OpenStudy (anonymous):

V=1/3Bh what is h expressed in terms of b and v

OpenStudy (xapproachesinfinity):

Just solve for h, \(\large V=1/3BH \Longrightarrow BH=3V \Longrightarrow H=\frac{3V}{B}\)

OpenStudy (anonymous):

thankyou so much

OpenStudy (xapproachesinfinity):

you are welcome!

OpenStudy (skullpatrol):

can you do the second part? @alexisdianna12